# Bilinear spherical maximal function with fractal dilations

### v1

@@ 5. Estimate for local maximal operator
$\bigg(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|L\left(U_{s,N}b_{1}\right)(x)\right|^{2}\;ds\bigg)^{\frac{1}{2}}$
$\leq 2^{n_{2}\frac{\beta}{2}}\int_{{\mathbb{R}}^{d}}\bigg(\int_{I}\left|U_{s,N}b_{1}(x-y)\right|^{2}\;ds\bigg)^{\frac{1}{2}}\frac{1}{\left(1+4\pi^{2}\left|y\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}dy$
$\leq 2^{n_{2}\frac{\beta}{2}}\int_{{\mathbb{R}}^{d}}\int_{{\mathbb{R}}^{d}}b_{1}(x-y-z)\Big(\int_{I}\left|K_{s,N}(z)\right|^{2}\;ds\Big)^{\frac{1}{2}}\frac{1}{\left(1+4\pi^{2}\left|y\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}~dzdy,$
@@
$\left\|\bigg(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|L\left(U_{s,N}b_{1}\right)\right|^{2}\;ds\bigg)^{\frac{1}{2}}\right\|_{L^{1}}\lesssim 2^{-n_{2}\frac{1-\beta}{2}}\|b_{1}\|_{L^{1}}.$
@@
$\Bigg\|\sum_{N=2^{n_{2}-1}}^{2^{n_{2}+1}}\mathcal{G}^{n_{2}}_{0,N}(b_{1},f_{2})\Bigg\|_{L^{2/3}}$ $\lesssim 2^{-n_{2}|\frac{2d-\beta}{2}}\bigg(\sum_{N=2^{n_{2}-1}}^{2^{n_{2}+1}}\|b_{1}\|_{L^{1}}^{\frac{2}{3}}\|F_{N}f_{2}\|_{L^{2}}^{\frac{2}{3}}\bigg)^{\frac{3}{2}}$
$\lesssim 2^{-n_{2}\frac{2d-\beta-2}{2}}\|b_{1}\|_{L^{1}}\|f_{2}\|_{L^{2}}.$
@@
$\Bigg\|\sum_{N=2^{n_{2}-1}}^{2^{n_{2}+1}}\tilde{\mathcal{G}}^{n_{2}}_{0,N}(b_{1},f_{2})\Bigg\|_{L^{2/3}}$ $\lesssim 2^{-n_{2}\frac{2d-\beta-4}{2}}\|b_{1}\|_{L^{1}}\|f_{2}\|_{L^{2}}.$
@@
$\big\|\mathcal{M}^{(0,n_{2})}_{0}(b_{1},f_{2})\big\|_{L^{2/3}}\lesssim 2^{-n_{2}\frac{2d-\beta-3}{2}}\|b_{1}\|_{L^{1}}\|f_{2}\|_{L^{2}}.$
This proves the desired bounds in (6.3) for $\mathcal{M}^{(0,n_{2})}_{0}$. The above argument can be modified suitably to get the desired bounds for $\mathcal{M}^{(0,n_{2})}_{k}$. ∎
Proof of part $(3)$ of Lemma 6.2.
@@
$|\{x\in\mathbb{R}^{d}:\mathcal{M}^{\mathbf{n}}(g_{1},g_{2})(x)>\frac{\alpha}{4}\}|$ $\lesssim$ $\frac{2^{-|\mathbf{n}|\frac{2d-2-\beta}{2}}\|g_{1}\|_{L^{2}}\|g_{2}\|_{L^{2}}\|_{L^{1}}}{\alpha}$
@@
We use similar arguments as in the previous section along with Lemma 6.4 and Lemma $5.2$ from [ChristZhou] to conclude that
@@
Interchanging the roles of the $i_{1}$ and $i_{2}$ proves the desired bound for $i_{2}$. This completes the proof of Lemma 6.4. Consequently, the proof of part $(3)$ of Lemma 6.2 is completed. ∎
@@
S.S. Choudhary is supported by the National Center for Theoretical Sciences under project no. 114-2124-M-002-004-. C.-Y. Shen is supported in part by NSTC through grant 111-2115-M-002-010-MY5.

### v2

@@ 5. Estimate for local maximal operator $\mathcal{M}_{E_{0}}$
$\bigg(\int_{I}\left|L\left(U_{s,N}b_{1}\right)(x)\right|^{2}\;ds\bigg)^{\frac{1}{2}}$ $\leq\int_{{\mathbb{R}}^{d}}\bigg(\int_{I}\left|U_{s,N}b_{1}(x-y)\right|^{2}\;ds\bigg)^{\frac{1}{2}}\frac{1}{\left(1+\left|y\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}dy$
$\leq\int_{{\mathbb{R}}^{d}}\int_{{\mathbb{R}}^{d}}b_{1}(x-y-z)\Big(\int_{I}\left|K_{s,N}(z)\right|^{2}\;ds\Big)^{\frac{1}{2}}\frac{1}{\left(1+\left|y\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}~dzdy,$
@@
$\left\|\bigg(\int_{I}\left|L\left(U_{s,N}b_{1}\right)\right|^{2}\;ds\bigg)^{\frac{1}{2}}\right\|_{L^{1}}\lesssim 2^{-\frac{n_{2}}{2}}\|b_{1}\|_{L^{1}}.$
@@
$\Bigg\|\sum_{N=2^{n_{2}-1}}^{2^{n_{2}+1}}\sum_{I\in\mathcal{E}_{0}}\mathcal{G}^{n_{2}}_{I,N}(b_{1},f_{2})\Bigg\|_{L^{2/3}}$ $\lesssim 2^{-n_{2}\frac{2d+1}{2}}\bigg(\sum_{N=2^{n_{2}-1}}^{2^{n_{2}+1}}\sum_{I\in\mathcal{E}_{0}}\|b_{1}\|_{L^{1}}^{\frac{2}{3}}\|F_{N}f_{2}\|_{L^{2}}^{\frac{2}{3}}\bigg)^{\frac{3}{2}}$
$\lesssim 2^{-n_{2}\frac{2d-1-3\beta}{2}}\|b_{1}\|_{L^{1}}\|f_{2}\|_{L^{2}}.$
@@
$\Bigg\|\sum_{N=2^{n_{2}-1}}^{2^{n_{2}+1}}\sum_{I\in\mathcal{E}_{0}}\tilde{\mathcal{G}}^{n_{2}}_{I,N}(b_{1},f_{2})\Bigg\|_{L^{2/3}}$ $\lesssim 2^{-n_{2}\frac{2d-3-3\beta}{2}}\|b_{1}\|_{L^{1}}\|f_{2}\|_{L^{2}}.$
@@
$\big\|\mathbb{I}^{n_{2}}\big\|_{L^{2/3}}\lesssim 2^{-n_{2}\frac{2d-2-3\beta}{2}}\|b_{1}\|_{L^{1}}\|f_{2}\|_{L^{2}}.$
Similarly, we can obtain the above estimate for $\mathbb{I}^{n_{2}}$. Now, combining the estimates for $\mathbb{I}^{n_{2}},\mathbb{II}^{n_{2}}$ and $\mathbb{III}^{n_{2}}$, we get that
$\big\|\mathcal{M}^{(0,n_{2})}_{0}(b_{1},f_{2})\big\|_{L^{2/3}}\lesssim 2^{-n_{2}\frac{2d-2-3\beta}{2}}\|b_{1}\|_{L^{1}}\|f_{2}\|_{L^{2}}.$
Note that $\max\left\{\frac{d-2-3\beta}{2},\frac{d-3-\beta}{2}\right\}\leq\frac{2d-2-3\beta}{2}$ for all $d\geq 3$ and $0\leq\beta\leq 1$. Therefore, the above estimate proves the desired bounds in (6.4) for $\mathcal{M}^{(0,n_{2})}_{0}$. ∎
6.3. Proof of Lemma 6.3.
@@
$|\{x\in\mathbb{R}^{d}:\mathcal{M}^{\mathbf{n}}(g_{1},g_{2})(x)>\frac{\alpha}{4}\}|$ $\lesssim\frac{2^{-|\mathbf{n}|\frac{2d-1-2\beta}{2}}\|g_{1}\|_{L^{2}}\|g_{2}\|_{L^{2}}}{\alpha}$
@@
We use similar arguments as in the previous section along with Lemma 6.5 and Lemma $5.2$ from [13] to conclude that
@@
Interchanging the roles of the $i_{1}$ and $i_{2}$ proves the desired bound for $i_{2}$. This completes the proof of Lemma 6.5. Consequently, the proof of Lemma 6.3 is completed. ∎
7. Proof of Theorem 2.2
First, observe that for any set $E\subseteq{\mathbb{R}}_{+}$, the maximal function $\mathcal{M}_{E}(f_{1},f_{2})$ is dominated in pointwise sense by the full maximal function $\mathcal{M}_{full}(f_{1},f_{2})$. Therefore, invoking the $L^{p}$-boundedness results of the operator $\mathcal{M}_{full}$, see [13, Theorem 1.3] and [16, Theorem 1], we get that the maximal operator $\mathcal{M}_{E}$ is bounded from $L^{p_{1}}\times L^{p_{2}}\to L^{p}$ for all $p_{1},p_{2}>2$.
Since,
$\mathcal{M}_{E}(f_{1},f_{2})(x)\leq\|f_{2}\|_{L^{\infty}}\sup_{t\in E}\left|\int_{\mathbb{S}^{1}}f_{1}(x+ty)~d\sigma(y,z)\right|.$
Observe that in order to study $L^{p}$ estimates of the maximal operator
$f_{1}\rightarrow\sup_{t\in E}\left|\int_{\mathbb{S}^{1}}f_{1}(x+ty)~d\sigma(y,z)\right|,$
it is enough to study the maximal operator
$\mathcal{N}f_{1}(x)=\sup_{t\in E}\int^{1}_{0}|f_{1}(x+tr)|~\frac{dr}{\sqrt{1-r^{2}}}.$
By Hölder’s inequality $\frac{1}{q}+\frac{1}{q^{\prime}}=1,~q\geq 1$, we have
$\mathcal{N}f_{1}(x)$ $\leq\sup_{t\in E}\sum_{j=1}^{\infty}2^{j\left(\frac{1}{2}-\frac{1}{q^{\prime}}\right)}\left(\int_{1-2^{-(j-1)}}^{1-2^{-j}}|f_{1}(x+tr)|^{q}~dr\right)^{\frac{1}{q}}$
$\leq\sum_{j=1}^{\infty}2^{-\frac{j}{2}}\sup_{k\in\mathbb{Z}}\sup_{t\in J_{k}}\left(\frac{1}{|I_{j}|}\int_{I_{j}}|f_{1}(x+2^{k}tr)|^{q}~dr\right)^{\frac{1}{q}},$
where $I_{j}=[1-2^{-(j-1)},1-2^{-j}]$ and $2^{k}J_{k}=E\cap[2^{k},2^{k+1}]$. Note that $J_{k}\subseteq[1,2].$ For each $k$, let $\mathcal{E}_{0,k}^{j}$ be the collection of disjoint intervals of length $2^{-j-1}$ that cover $J_{k}$. Then, Definition 1.1 tells us that $\sup_{k\in{\mathbb{Z}}}\#\mathcal{E}_{0,k}^{j}\leq 2^{j\beta}$. By a change of variable $tr\to r$, we have
$\sup_{t\in J_{k}}\left(\frac{1}{|I_{j}|}\int_{I_{j}}|f_{1}(x+2^{k}tr)|^{q}~dr\right)^{\frac{1}{q}}$ $=\sup_{J\in\mathcal{E}_{0,k}^{j}}\sup_{t\in J}\left(\frac{1}{t|I_{j}|}\int_{tI_{j}}|f_{1}(x+2^{k}r)|^{q}~dr\right)^{\frac{1}{q}}$
$\lesssim\left(\sum_{J\in\mathcal{E}_{0,k}^{j}}\frac{1}{|\tilde{I_{j}}|}\int_{\tilde{I_{j}}}|f_{1}(x+2^{k}r)|^{q}~dr\right)^{\frac{1}{q}},$
where $\tilde{I_{j}}=\{ab:a\in J,b\in I_{j}\}$ with $|\tilde{I_{j}}|\approx|I_{j}|$, and we used $\ell_{q}\hookrightarrow\ell_{\infty}$ for the set $\mathcal{E}_{0,k}^{j}$ in the last inequality. Therefore,
$\mathcal{N}f_{1}(x)\lesssim\sum_{j=1}^{\infty}2^{-j\left(\frac{1}{2}-\frac{\beta}{q}\right)}\sup_{k\in\mathbb{Z}}\left(\frac{1}{|\tilde{I_{j}}|}\int_{\tilde{I_{j}}}|f_{1}(x+2^{k}r)|^{q}~dr\right)^{\frac{1}{q}}.$
Thus, for $p>q$, we have
$\left\|\mathcal{N}f_{1}\right\|_{L^{p}}$ $\lesssim\sum_{j=1}^{\infty}2^{-j\left(\frac{1}{2}-\frac{\beta}{q}\right)}\left\|\sup_{k\in\mathbb{Z}}\left(\frac{1}{|\tilde{I_{j}}|}\int_{\tilde{I_{j}}}|f_{1}(x+2^{k}r)|^{q}~dr\right)^{\frac{1}{q}}\right\|_{L^{p}}$
$\lesssim\sum_{j=1}^{\infty}2^{-j\left(\frac{1}{2}-\frac{\beta}{q}\right)}\log(2+c2^{j})\|f_{1}\|_{L^{p}},$
where we have used the $L^{p}$-estimates of shifted Hardy-Littlewood maximal function, see [31, Theorem 4.1] and [39] for more details. The sum in the above inequality is finite if $q>2\beta$. This implies that $\mathcal{M}_{E}$ is bounded from $L^{p}({\mathbb{R}})\times L^{\infty}({\mathbb{R}})$ to $L^{p}({\mathbb{R}})$ for $p>\max\{2\beta,1\}$. Interchanging the role of $f_{1}$ and $f_{2}$, we can obtain $L^{\infty}\times L^{p}\to L^{p}$-boundedness of $\mathcal{M}_{E}$ for $p>\max\{2\beta,1\}$.
Interpolating the above estimates with trivial estimates from $\mathcal{M}_{full}$, we get the desired boundedness in Theorem 2.2 when $\beta\geq\frac{1}{2}$. Moreover, if $\beta<\frac{1}{2}$, $\mathcal{M}_{E}$ maps $L^{p_{1}}({\mathbb{R}})\times L^{p_{2}}({\mathbb{R}})$ to $L^{p}({\mathbb{R}})$, for $1<p_{1},p_{2}<\infty$.
Next, following Littlewood-Paley decomposition from (3), we have
$\mathcal{M}_{E}(f_{1},f_{2})(x)\lesssim$ $\sup_{k\in{\mathbb{Z}}}\sup_{t\in E_{k}}\Big|\mathcal{A}_{t}(P_{k}f_{1},P_{k}f_{2})(x)\Big|+\sup_{k\in{\mathbb{Z}}}\sup_{t\in E_{k}}\Big|\mathcal{A}_{t}(P_{k}f_{1},f_{2})(x)\Big|$ (7.1)
$+\sup_{k\in{\mathbb{Z}}}\sup_{t\in E_{k}}\Big|\mathcal{A}_{t}(f_{1},P_{k}f_{2})(x)\Big|+\sum^{\infty}_{n_{1},n_{2}=1}\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})(x),$
where
$\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})(x)=\sup_{k\in{\mathbb{Z}}}\sup_{t\in E_{k}}\Big|\mathcal{A}_{t}(R_{k+n_{1}}f_{1},R_{k+n_{2}}f_{2})(x)\Big|.$
Recall from the proof of Lemma 3.1 that
$\sup_{t\in E_{k}}\sup_{|z|\leq 1}\big|P_{k}f_{2}(x-tz)\big|\lesssim M_{HL}f_{2}(x).$
Then, we have
$\left\|\sup_{k\in{\mathbb{Z}}}\sup_{t\in E_{k}}\Big|\mathcal{A}_{t}(f_{1},P_{k}f_{2})\Big|\right\|_{L^{p}}$ $\lesssim\big\|M_{HL}f_{2}\big\|_{L^{p_{2}}}\left\|\sup_{t\in E}\left|\int_{\mathbb{S}^{1}}f_{1}(\cdot+ty)~d\sigma(y,z)\right|\right\|_{L^{p_{1}}}.$
From the boundedness of the Hardy-Littlewood maximal function and the boundedness of the operator $\sup_{t\in E}\left|\int_{\mathbb{S}^{1}}f_{1}(x+ty)~d\sigma(y,z)\right|$ discussed before, the maximal operator $\sup_{k\in{\mathbb{Z}}}\sup_{t\in E_{k}}\big|\mathcal{A}_{t}(f_{1},P_{k}f_{2})\big|$ is bounded from $L^{p_{1}}\times L^{p_{2}}$ to $L^{p}$ for $p_{1},p_{2}>1$ when $\beta<\frac{1}{2}$. Similarly, we can get the boundedness of other two terms in (7.1) involving $P_{k}$.
Thus, it suffices to prove that for each $\mathbf{n}$, the maximal operator $\mathcal{M}^{\mathbf{n}}$ maps $L^{p_{1}}\times L^{p_{2}}$ into $L^{p}$ with bound $O\big(2^{-\epsilon|\mathbf{n}|}\big)$ for some $\epsilon>0$ in desired range of boundedness in Theorem 2.2. Since $\beta<\frac{1}{2}$, using the boundedness of $\mathcal{M}_{E}$, we have
$\left\|\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})\right\|_{L^{p}}\lesssim\|f_{1}\|_{L^{p_{1}}}\|f_{2}\|_{L^{p_{2}}},$
when $1<p_{1},p_{2}<\infty$. Interpolating the estimates with Lemma 6.1 and Lemma 6.3, we get
$\left\|\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})\right\|_{L^{p}}\lesssim 2^{-\epsilon|\mathbf{n}|}\|f_{1}\|_{L^{p_{1}}}\|f_{2}\|_{L^{p_{2}}},$
for some $\epsilon>0$ when $\left(\frac{1}{p_{1}},\frac{1}{p_{2}}\right)\in\Omega(O,D_{1},D_{2},D_{3})$. Now, summing in $\mathbf{n}$ completes the proof of Theorem 2.2.
7.1. Example.
Let $E\subseteq[1,2]$ with Minkowski dimension $\beta$. For a small $\delta>0$, let $\{I_{j}\}_{j=1}^{\delta^{-\frac{\beta}{2}}}$ be a collection of $\delta^{\frac{1}{2}}$-length disjoint intervals that cover $E$. Define the set $F=\cup_{j=1}^{\delta^{-\frac{\beta}{2}}}4I_{j}$, where $4I_{j}$ is the interval of length $4\delta^{\frac{1}{2}}$ with the same center as $I_{j}$. Then $|F|\lesssim\delta^{\frac{1-\beta}{2}}$.
Let $f_{1}(x)={\chi}_{[-\delta,\delta]}$ and $f_{2}(x)={\chi}_{F}(x)$. Then, for $x$ in $\delta$-neighbourhood of $E$, we have
$\mathcal{M}_{E}(f_{1},f_{2})(x)\gtrsim\delta^{\frac{1}{2}}.$
Thus, if $\mathcal{M}_{E}$ is bounded from $L^{p_{1}}\times L^{p_{2}}$ to $L^{p}$, then
$\delta^{\frac{1}{2}}\delta^{\frac{1-\beta}{p}}\lesssim\|\mathcal{M}_{E}(f_{1},f_{2})\|_{L^{p}}\lesssim\|f\|_{L^{p_{1}}}\|f_{2}\|_{L^{p_{2}}}=\delta^{\frac{1}{p_{1}}}\delta^{\frac{1-\beta}{2p_{2}}}.$
Taking the limit $\delta\to 0$, we get that
$\frac{1}{p_{1}}+\frac{1-\beta}{2p_{2}}\leq\frac{1-\beta}{p}+\frac{1}{2}.$
Interchanging the role of functions $f_{1}$ and $f_{2}$, we can get that
$\frac{1-\beta}{2p_{1}}+\frac{1}{p_{2}}\leq\frac{1-\beta}{p}+\frac{1}{2}.$
Remark 7.1.
If we assume $\beta\geq\frac{1}{2}$ in the above equations along with either $p_{2}=\infty$ or $p_{1}=\infty$, then We get that $p_{1}\geq 2\beta$ or $p_{2}\geq 2\beta$. Note that from our result we recover sharp boundedness for this case except the endpoint.
@@
Surjeet Singh Choudhary is supported by the National Center for Theoretical Sciences through NSTC grant 114-2124-M-002-004. Chun Yen Shen is supported in part by NSTC through grant 111-2115-M-002-010-MY5. Saurabh Shrivastava acknowledges the support from Anusandhan National Research Foundation (ANRF), India under the project ANRF/ARG/2025/000940/MS.